OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. solve |4x+3|=9+2x

OpenStudy (anonymous):

I think its x=3 and x=-3 but im want to double check

OpenStudy (e.cociuba):

Separate it into 2 equations when there is an absolute value: Solve: 4x+3 = 9+2x and 4x+3 = -(9+2x)

OpenStudy (e.cociuba):

Get x by itself by combining like terms.

OpenStudy (e.cociuba):

One of your answers are correct, the other is wrong.

OpenStudy (anonymous):

when I solved it that way I got x=3 and x=-2

OpenStudy (e.cociuba):

Correct :)

OpenStudy (anonymous):

yayy! what about extraneous solutions? are there any? I have no clue what those are

OpenStudy (e.cociuba):

I believe there are none. Not quite sure though. It's when solutions you have algebraically derived, but do not conform to the stated limitations or hypotheses of the original problem.

OpenStudy (anonymous):

okay I think we got it thank you so much!!

OpenStudy (e.cociuba):

The question is not asking for it, so I would leave it at what you have. It just say's "check" for them :)

OpenStudy (e.cociuba):

No problemo! :)